OpenStudy (anonymous):

An airplane flew 3.5 hours with a 40 mph head wind. The return trip with a tail wind of the same speed took 2.5 hours. Find the speed of the plane in still air. Answer 220 mph 200 mph 180 mph 240 mph An airplane flew 3.5 hours with a 40 mph head wind. The return trip with a tail wind of the same speed took 2.5 hours. Find the speed of the plane in still air. Answer 220 mph 200 mph 180 mph 240 mph @Mathematics

hero (hero):

Find the distance first

hero (hero):

d = rt

hero (hero):

d = rt d = (x - 40)*3.5 d = (x + 40)*2.5 d = d 3.5x - 140 = 2.5x +100 x = 240 The plane is flying 240mph